OpenStudy (anonymous):

The slope of any line perpendicular to , where A(0, -3) and B(4, 9) are two points on the line, is m =

OpenStudy (zzr0ck3r):

find slope with the two points slope = (9-(-3))/(4-0) = 12/4 = 3 so when the slope of a line is three, the slope of the line orthogonal to that line is the negative inverse of that slope so -1/3